# Generating coupled cluster code for modern distributed memory tensor software

### v1

@@ Abstract
Scientific groups are struggling to adapt their codes to quickly-developing GPU-based HPC platforms. The domain of distributed coupled cluster (CC) calculations is not an exception. Moreover, our applications to tiny QED effects require higher-order CC which include thousands of tensor contractions, which makes automatic treatment imperative. The challenge is to allow efficient implementation by capturing key symmetries of the problem, while retaining the abstraction from the hardware. We present the tensor programming framework tenpi, which seeks to find this balance. It features a python library user interface, global optimization of intermediates, a visualization module and Fortran code generator that bridges the DIRAC package for relativistic molecular calculations to tensor contraction libraries. tenpi brings higher-order CC functionality to the massively parallel module of DIRAC. The architecture and design decision schemes are accompanied by benchmarks and by first production calculations on Summit, Frontier and LUMI along with state-of-the-art of tensor contraction software.
@@ I INTRODUCTION
According to Jack Dongarra, the cofounder of the TOP500 list, 99% of the FLOP performance of modern supercomputers lies in GPU accelerators.Dongarra (2023); Chirigati (2022) However, in the domain of distributed coupled cluster calculations, which is the focus of the present work, most implementations are able to use only about 10% of the theoretical maximum FLOP rate.Calvin et al. (2021)
This exposes the difficult situation in which scientists find themselves.Ma, Krishnamoorthy, and Agrawal (2011); Woolston (2022) Writing down fixed BLAS and MPI statements is no longer sufficient as modern machines have heterogeneous structures.Lyakh (2019) Figure 1 contains a schematic representation of a node of the Summit supercomputer. Communication throughputs and memory sizes differ substantially within and among the machines. Restructuring data movement across multiple levels of hierarchy to reduce the cost is a challenging problem,Ma, Krishnamoorthy, and Agrawal (2011) where the developer basically faces a graph theoretical task.Thibault (2018) The traditional 5-year lifespan of supercomputersRojas et al. (2019) has accelerated to about 3 years due to AI-driven breakthroughs in energy efficiencyZhao et al. (2023) and keeping up with the changes requires expert manpower for which academia competes with industry.Woolston (2022); Heffernan and McKay (2018)
This develops a pressure on systematic treatment of the development.Ma, Krishnamoorthy, and Agrawal (2011); Lechner et al. (2024); Papadopoulos (2019); Auer et al. (2006) One way is to build the code such that it can be quickly adapted to underlying numerical software changes using a code generator.Lechner et al. (2024); Papadopoulos (2019); Auer et al. (2006) Another way is to rely on tools which allow a high level of abstraction, like MATLAB,The Mathworks, Inc. (2024) MapleWaterloo Maple, Inc. (2024) or Mathematica,Inc. (2024) to write in parallel frameworks like SYCLThe Khronos Group, Inc. (2021) and OpenACC,The OpenACC Organization (2022) to use runtime environments like StarPU,Augonnet et al. (2011) MADNESSHarrison et al. (2016) and LEGION,Bauer et al. (2012) or tensor compilers like DISTAL.Yadav, Aiken, and Kjolstad (2022)
@@
The challenge is to allow efficient implementation by capturing key mathematical symmetries of the problem, while retaining the abstraction from the hardware.Matthews (2018a) Strong candidates for finding this balance are tensor compilersYadav, Aiken, and Kjolstad (2022) which have seen an explosion of interestZheng et al. (2024) in the computer science (CS) community. However, to enable production use by chemists, performance-critical features like index permutation symmetry or block sparsity would need to be implemented.Panyala (2014); Shiozaki et al. (2008) The same holds for major machine learning libraries. Unlike in the case of AI, where big tech companies drive the development, in coupled cluster and tensor networks, scientists are left to rely on themselves. This was concluded by the representatives of big technology companies at the CECAM workshopBrandejs et al. (2024) which we organized on the topic.
The purpose of this paper is to explore the path from the high-level representation of CC methods to an efficient code on modern machines and to provide the reader with a state-of-the art on the necessary tensor software.
@@ I.1 Elegant development of CC methods, systematic approach
Equations in coupled cluster theory can reach a degree of complexity where manual manipulation is no longer practical. Higher orders of the method with quadruple excitations and above required for accurate applications include hundreds of terms with many indices. Even though derivation by hand can be simplified by using diagrams which help detect equivalent terms, further manipulations like the design of intermediates with respect to a given cost-functionLai et al. (2012), implementing other flavors of the method or verifying the correctness of the derivation steps are very error-prone when done by hand and hence are clear candidates for automatic symbolic treatment, which can speed up the implementation significantly.
There have been numerous efforts in this direction with differing degrees of success. Here we focus on complete toolchains which include both equation and code generation - these are often linked to an existing quantum chemistry package. A prominent example and one of the pioneering systematization efforts is the Tensor Contraction Engine (TCE)Baumgartner et al. (2005) initially developed by Hirata,Auer et al. (2006) a part of NWCHEM software.Kendall et al. (2000) Even though TCE was not the such code generator,Janssen and Schaefer (1991); Li and Paldus (1994); Kállay and Surján (2001); Nooijen and Lotrich (2001) to the best of our knowledge, it has the most complete description in literature.Lechner et al. (2024) Numerous works follow its scheme:Shiozaki et al. (2008)
@@
Another noteworthy abstraction effort is the domain specific language (DSL) for tensor operations SIAL (superinstruction assembly language) in the ACES III package.Lotrich et al. (2009); Deumens et al. (2011) Even though initially successful, its limitation was the same as that of TCE: customizability. Not all CC methods can be expressed in the form of tensor contraction expressions and without C++ or Fortran frontend support, one could not easily include custom code and optimizations.Panyala (2014) Further packages using their own tensor DSL include QChemShao et al. (2006) based on the libtensorEpifanovsky et al. (2013) library, and Cyclops CTFSolomonik et al. (2013) whose performance for distributed CPU contractions gained fame in the computer science community, where it is used as a reference.Calvin, Lewis, and Valeev (2015); Yadav, Aiken, and Kjolstad (2022); Irmler et al. (2023)
Formula generators themselves divide in three main groups based on their approach to the derivation: by applying the algebra of creation and annihilation operators, by Wick’s theorem or diagrams.
@@
Recent ORCANeese (2022) is good example of a large-scale systematization effort where a substantial part of the package is generated using the ORCA-AGE generatorLechner et al. (2024) based directly on application of the algebra of creation and annihilation operators. Even though this approach is very general, it took up to 7 years to resolve performance issues with the generation.Krupička et al. (2017); Lechner et al. (2024)
New efforts have emerged in past years, as groups behind quantum chemical packages consider having their own toolchain. Most of them use Wick’s theorem: SeQuant of the Valeev group Bochevarov, Valeev, and David SheRrill (2004) which is a C++ rewrite of their Mathematica code, Drudge/Gristmill of the Scuseria groupSong, Henderson, and Scuseria (2022) which can switch between different abstract algebras and where the use of Wick’s theorem is optional, FEMTO of SaitowSaitow, Kurashige, and Yanai (2013) used for DMRG-MRCI and pair-natural orbitals, SQANeuscamman, Yanai, and Chan (2009) and GeCCo.Köhn, Richings, and Tew (2008) These efforts can be further classified by supported methods and features, see more detailed reviews.Lechner et al. (2024); Solomonik et al. (2014) Many other efforts exist which focus on only one of the three steps of the process.G. A. Smith and Gray (2018); Quintero-Monsebaiz and Loos (2023); Lyakh, Ivanov, and Adamowicz (2004); Li and Paldus (1994)
@@ I.2 Optimization step
The most difficult part of the entire process is the design of intermediates. This problem has not only proven to be NP-hard,Chi-Chung, Sadayappan, and Wenger (1997); Lechner et al. (2024) but in practice the actual cost-function depends on the model system, the hardware and the underlying math library. Instead, most optimizers rely on a naive FLOP count cost model, despite the fact that in the communication-heavy distributed CC calculations, this is not even the leading term. Most chemistry packages use just basic heuristics of single-term contraction path optimization and do not implement multi-term optimization for which they rely on hand-tuning.
The TCE also relies on performance model-driven search-based approach to program transformation, using a global optimizer OpMin written by Sadayappan.Lai et al. (2012) The optimization consists of:
@@
which later became a typical structure. OpMin performs well compared with hand-tuned code of NWCHEM for higher-order CC also thanks to its index-permutation-symmetry-awareness. But it does not support perturbative approaches with energy denominators.
Other efforts include a distributed GPU-contraction optimizerMa, Krishnamoorthy, and Agrawal (2011) benchmarked for CCSD(T), followed by the AutoHOOT optimizerMa, Ye, and Solomonik (2020) (not CC), and an optimizer with a GPU-aware cost model.Harju et al. (2013) The aforementioned complete toolchains also include own optimizers of varying levels of sophistication. For more information, see the literature overview in the thesis of PanyalaPanyala (2014) on loop-level optimization.Gao et al. (2006); Sahoo et al. (2005) Note that there is no clear boundary on how low-level an optimizer should be. If it reaches too low-level, it becomes a compiler, which is in fact a trend in state-of-the-art works.Yadav, Aiken, and Kjolstad (2022); Yadav et al. (2024) Despite a strong connection to compiler optimizations for different applications,Johnson et al. (2001); Meirom et al. (2022); Bondhugula et al. (2008) these cannot be directly applied to tensor contractions in CC due to their distinctive featuresPanyala (2014):
@@
A specific index permutation symmetry
@@ I.3 State-of-the-art: Tensor contraction
Tensor contraction represents the most computer-intensive operation of numerous methods in quantum chemistry, condensed matter physics, nuclear physics, machine learning and quantum computing.Psarras et al. (2021); Matthews (2018b); Brabec et al. (2020) Examples include coupled cluster methods,Calvin et al. (2021); Peng et al. (2016); Herault et al. (2021) tensor networks,Levy, Solomonik, and Clark (2020) quantum computing simulatorsLyakh et al. (2022) and certain neural networks Sharir, Shashua, and Carleo (2022) and signal processing methods.Chi and Kolda (2012); Sidiropoulos et al. (2017) There, the key limitation to the system size is the cost of tensor contraction.
Typically, the software packages decompose tensor contractions into primitive matrix operations Shi et al. (2016); Georganas et al. (2021) and pass them to BLAS. There are different approaches of dealing with tensor transposition (reshape), which is in general required for the usage of GEMM, which depends on a unit-stride index to multiply two matrices. (Note that in this context, the CS community uses the term transposition interchangeably with reshape.) The most common approach is TTGTMatthews (2018b) (Transpose-Transpose-GEMM-Transpose): for
@@
In his block-scatter matrix tensor contraction (BSMTC) scheme, MatthewsMatthews (2018b) has shown that is possible to avoid transposition entirely by loading parts of tensors into CPU L2 and L3 cache in a specific order.
Another alternative is the GETT scheme (GEMM-like tensor-tensor multiplication) by Springer and BientinesiSpringer and Bientinesi (2016) which generates code to call matrix-matrix multiplication kernels while again reformulating the sub-matrix-packing. A GPU implementation thereof has become a foundation of the proprietary single-node library cuTENSOR.cuT (2019)
@@
Compared to the software for matrix operations, we have identified two major practical hurdles for tensor software:
@@
Contrary to BLAS for matrix operations, there is no standardized interface for tensor operations.Psarras et al. (2021) Currently, there is an ongoing standardization push attempting to find agreement across academia and companies.
@@
As for standardization, there has been some work in the past, but none of the interfaces have prevailed. BTASShi et al. (2016) has aimed at providing a standard and a basic CPU implementation. TBLISMatthews (2018b) provides BLAS-like tensor calls. Until today, there is no GPU implementation of these. During the Dagstuhl tensor workshop Bientinesi et al. (2022) in 2022, a development of domain-specific tensor language has been initiated, but remains unfinished and undocumented. There has been work-technical specifications of additional aspects of standardization, like tensor-memory distributionValeev (2023); cppreference.com community (2023) and randomized multilinear algebra.Murray et al. (2023)
Regarding the implementations, there are numerous scattered efforts listed in Ref. Psarras et al., 2021. However, for open-source distributed memory GPU libraries, there is only a few major players. i) ExaTENSORLyakh (2019); Pototschnig et al. (2021) based on TAL-SHtal (2018), where the key developer has left to industry. ii) TiledArrayPeng et al. (2016) and Cyclops CTF,Solomonik et al. (2013) where the solid support for GPUs is still under development. iii) TACO DISTAL,Yadav, Aiken, and Kjolstad (2022) which has performance issues for higher-order tensors. iv) TAMMMutlu et al. (2023) scales well on supercomputers, but has a potentially problematic dependency of Global Arrays,Palmer et al. (2022) with a small user base for a communication library. v) cuTENSORZhang et al. (2020) is proprietary and only supports Nvidia hardware, while vi) hipTensorMillette, Ma, and Karunanidhi (2024) is at an early stage of development. Other industry products like vii) PyTorchPaszke et al. (2019) and viii) TensorFlowAbadi (2016) focus almost exclusively on machine learning and lack features required by other application domains.

### v2

@@ Abstract
Using GPU-based HPC platforms efficiently for coupled cluster computations is a challenge due to heterogeneous hardware structures. The constant need to adapt software to these structures and the required man-hours makes a systematization of high-performance code development desirable, even more so for higher-order coupled cluster. This is generally achieved by introducing a high-level representation of the problem, which is then translated to low-level instructions for the hardware using a compiler/translator component. Designing such software comes with another challenge: Allowing efficient implementation by capturing key symmetries of tensors, while retaining the abstraction from the hardware. We review ways to address these two challenges while presenting design decisions which led us to the development of a general-order coupled cluster code generator. The systematically produced code shows excellent weak scaling behavior running on up to 1200 GPUs using the distributed memory tensor library ExaTENSOR. We present an open-source modular tensor framework "tenpi" for coupled cluster code development with diagrammatic derivation, visualization module, symbolic algebra, intermediate optimization and support for multiple tensor backends. Tenpi brings higher-order CC functionality to the massively parallel ExaCorr module of the DIRAC code for relativistic molecular calculations.
@@ I INTRODUCTION
According to Jack Dongarra, the cofounder of the TOP500 list,TOP500.org (2024) 99% of the FLOP performance of modern supercomputers lies in GPU accelerators.Dongarra (2023); Chirigati (2022) However, in the domain of distributed coupled cluster calculations, the focus of the present work, most implementations are able to use only about 10% of the theoretical maximum FLOP rate.Calvin et al. (2021)
This exposes the difficult situation in which scientists find themselves.Ma, Krishnamoorthy, and Agrawal (2011); Woolston (2022) Implementing fixed BLAS (Basic Linear Algebra Subroutines) and MPI (Message-Passing Interface) statements is no longer sufficient as modern machines have heterogeneous structures.Lyakh (2019) Fig. 1 contains a schematic representation of a node of the Summit supercomputer. Communication throughputs and memory sizes differ between the components by orders of magnitude. Restructuring data movement across multiple levels of hierarchy to reduce the cost is a challenging problem,Ma, Krishnamoorthy, and Agrawal (2011) where the developer basically faces a graph theoretical task.Thibault (2018) The traditional 5-year lifespan of supercomputersRojas et al. (2019) has accelerated to about 3 years due to AI-driven breakthroughs in energy efficiencyZhao et al. (2023) and keeping up with the changes requires expert manpower for which academia competes with industry.Woolston (2022); Heffernan and McKay (2018)
The situation benefits systematic approaches to development and parallelization.Ma, Krishnamoorthy, and Agrawal (2011); Lechner et al. (2024); Papadopoulos (2019); Auer et al. (2006) One way is to build the code such that it can be quickly adapted to underlying numerical software changes using a code generator.Lechner et al. (2024); Papadopoulos (2019); Auer et al. (2006) Another way is to rely on tools which allow a high level of abstraction, like MATLAB,The Mathworks, Inc. (2024) MapleWaterloo Maple, Inc. (2024) or Mathematica,Wolfram Research, Inc. (2024) to write in parallel frameworks like SYCLThe Khronos Group, Inc. (2021) and OpenACC,The OpenACC Organization (2022) to use runtime environments like StarPU,Augonnet et al. (2011) PaRSEC,Herault et al. (2021) MADNESS,Harrison et al. (2016) and LEGION,Bauer et al. (2012) or tensor compilers like DISTAL.Yadav, Aiken, and Kjolstad (2022)
@@
The challenge is to allow efficient implementation by capturing key mathematical features and symmetries of the problem, while retaining the abstraction from the hardware.Matthews (2018a) Strong candidates for finding this balance are tensor compilersYadav, Aiken, and Kjolstad (2022) which have seen an explosion of interestZheng et al. (2024) in the computer science (CS) community. However, to enable production use in the coupled cluster domain, performance-critical features like index permutation symmetry or block sparsity would need to be implemented.Panyala (2014); Shiozaki et al. (2008) The same holds for major machine learning libraries. Unlike in the case of AI, where big tech companies drive the development, in coupled cluster and tensor networks, scientists are left to rely on themselves. This was concluded at a panel discussion with representatives of these companies at a recent CECAM workshopBrandejs et al. (2024) which we co-organized.
The purpose of this paper is to explore the path from a high-level representation of CC methods to an efficient code on modern machines from the viewpoint of software-architecture. We provide the reader with a state-of-the art on the necessary tensor software, and we benchmark tenpiBrandejs and Saue (2024) - our solution to the two challenges: i) adapting CC to heterogeneous HPC, ii) identifying a balanced high-level representation.
@@ I.1 Elegant development of CC methods, systematic approach
Equations in coupled cluster theory can reach a degree of complexity where manual manipulation is no longer practical.Kállay and Surján (2001) Higher orders of the method with quadruple excitations and above include hundreds of terms with many indices. This accuracy is required to reach the scale of quantum-electrodynamics (QED) effects which our group is invested in.Saue et al. (2024)
Even though derivation by hand can be simplified by using diagrams which help detect equivalent terms, subsequent work with many indices is error-prone when done by hand. This and further manipulations like the design of intermediates with respect to a given cost-functionLai et al. (2012) or implementing other flavors of the method are clear candidates for automatic symbolic treatment.
There have been numerous efforts in this direction with differing degrees of success. Here we focus on complete toolchains which include both equation and code generation - these are often linked to an existing quantum chemistry package. A prominent example and one of the pioneering systematization projects is the Tensor Contraction Engine (TCE)Baumgartner et al. (2005) initially developed by Hirata,Auer et al. (2006) a part of NWCHEM software.Kendall et al. (2000) Even though TCE was not the first such code generator,Janssen and Schaefer (1991); Li and Paldus (1994); Kállay and Surján (2001); Nooijen and Lotrich (2001) to the best of our knowledge, it has the most complete description in literature.Lechner et al. (2024) Later works often follow its scheme:Shiozaki et al. (2008)
@@
Another noteworthy abstraction effort is the domain-specific language (DSL) SIAL (superinstruction assembly language) for tensor operations in the ACES III package.Lotrich et al. (2009); Deumens et al. (2011) Even though initially successful, its limitation was the same as that of TCE: customizability. Not all methods from the CC family can be efficiently expressed in the form of tensor contractions and without C++ or Fortran frontend support, one could not easily include custom code and optimizations.Panyala (2014) Some key parallelization optimizations even had to be done by hand in final Fortran loops of the generated code, as it was too difficult to modify the generator itself.Ozog et al. (2013) Further packages using their own tensor DSL include QChemShao et al. (2006) based on the libtensorEpifanovsky et al. (2013) library, and Cyclops CTFSolomonik et al. (2013) whose performance for distributed CPU contractions gained distinction in the computer science community, where it is used as a reference.Calvin, Lewis, and Valeev (2015); Yadav, Aiken, and Kjolstad (2022); Irmler et al. (2023)
CC formula generators themselves divide in three main groups based on their approach to the derivation: i) by applying the algebra of creation and annihilation operators, ii) by Wick’s theorem, iii) or by diagrams.
@@
Recent ORCANeese (2022) is good example of a large-scale systematization effort where a substantial part of the package is generated using the ORCA-AGE generatorLechner et al. (2024) based directly on application of the algebra of creation and annihilation operators. Even though this approach is very general, it took up to 7 years before performance issues with the generation were resolved.Krupička et al. (2017); Lechner et al. (2024)
New efforts have emerged in past years, as groups behind quantum chemical packages consider having their own toolchain. Most of them use Wick’s theorem: i) SeQuant of the Valeev group Bochevarov, Valeev, and David Sherrill (2004) which is a C++ rewrite of their Mathematica code, ii) Drudge/Gristmill of the Scuseria groupSong, Henderson, and Scuseria (2022) which can switch between different abstract algebras and where the use of Wick’s theorem is optional, iii) FEMTO of SaitowSaitow, Kurashige, and Yanai (2013) used for DMRG-MRCI and pair-natural orbitals, iv) SQANeuscamman, Yanai, and Chan (2009), v) ebccBackhouse (2024) and vi) GeCCo.Köhn, Richings, and Tew (2008) These efforts can be further classified by supported methods and features, see more detailed reviews.Lechner et al. (2024); Solomonik et al. (2014) Many other efforts exist which focus on only one of the three steps of the process.G. A. Smith and Gray (2018); Quintero-Monsebaiz and Loos (2023); Lyakh, Ivanov, and Adamowicz (2004); Li and Paldus (1994); Wang, Peyton, and Crawford (2022)
@@ I.2 Optimization step
Algorithmically the most complicated part of the entire process is the automatic design of intermediates. Finding the optimal set of intermediates is NP-hard,Chi-Chung, Sadayappan, and Wenger (1997); Lechner et al. (2024) so heuristics is used in practice. The ideal cost function would be the actual walltime of the target calculation, which depends on the size of the application system, the hardware and the underlying math library. As this is unknown a priori, most optimizers rely on a naive FLOP count cost model, despite the fact that distributed CC calculations are often communication-bound, meaning that this model does not take into account the leading term of the walltime. Most chemistry codes are limited only to the basic single-term contraction path optimization and do not implement global (multi-term) optimization, which is replaced by hand-tuning.
The TCE also relies on search-based approach driven by a FLOP count performance model, using a global optimizer OpMin written by Sadayappan et al.Lai et al. (2012) The optimization consists of:
@@
This is a common structure among tensor contraction optimizers, with differing level of sophistication.Panyala (2014) OpMin performs well compared with hand-tuned code of NWCHEM for higher-order CC, also because its cost-function takes into account the index-permutation symmetry. Nevertheless, it does not support perturbative approaches with energy denominators.
Other efforts include a distributed GPU-contraction optimizer benchmarked for CCSD(T),Ma, Krishnamoorthy, and Agrawal (2011) followed by the AutoHOOT optimizerMa, Ye, and Solomonik (2020) (not CC), and an optimizer with a GPU-aware cost model.Harju et al. (2013) The aforementioned complete toolchains also include own optimizers of varying levels of sophistication. For more information, see the literature overview in the thesis of PanyalaPanyala (2014) on loop-level optimization.Gao et al. (2006); Sahoo et al. (2005) Note that there is no clear boundary on how low-level an optimizer should be. If it reaches low-level, it can be called a compiler, which is in fact a trend in state-of-the-art works.Yadav, Aiken, and Kjolstad (2022); Yadav et al. (2024); Zheng et al. (2024) There is a strong connection between tensor compilers and compiler optimizations for different applications.Johnson et al. (2001); Meirom et al. (2022); Bondhugula et al. (2008) However, CC contractions require specific treatment due to their distinctive features:Panyala (2014)
@@
Dependencies preventing loop fusion never occur
Last but not least, a noteworthy effort is the load balancer DLTCLai et al. (2013) which analyzes dependencies between contractions and groups them in layers that can be executed concurrently. This is particularly relevant for higher-order CC with large number of contractions with widely different computational cost.Solomonik et al. (2014)
@@ I.3 State-of-the-art: Tensor contraction
Tensor contraction represents the most computer-intensive operation of numerous methods in quantum chemistry, condensed matter physics, nuclear physics, machine learning and quantum computing.Psarras et al. (2022); Matthews (2018b); Brabec et al. (2020) Examples include coupled cluster methods,Calvin et al. (2021); Peng et al. (2016); Herault et al. (2021) tensor networks,Levy, Solomonik, and Clark (2020) quantum computing simulators,Lyakh et al. (2022) certain neural networks Sharir, Shashua, and Carleo (2022) and signal processing methods.Chi and Kolda (2012); Sidiropoulos et al. (2017) Common to all these is that the key limitation to the affordable system size is the cost of tensor contraction.
Typically, the software packages decompose tensor contractions into primitive matrix operations and pass them to BLAS.Shi et al. (2016); Georganas et al. (2021) There are different approaches of dealing with tensor transposition (reshape), which is in general required for the usage of BLAS GEMM (General Matrix Multiplication), which depends on a unit-stride index to multiply two matrices. Given a tensor memory layout, an index is said to have a unit stride when incrementing it by one translates to a move by one scalar in memory. Note that in the context of tensor operations, the CS community uses the term transposition interchangeably with reshape. The most common approach is TTGTMatthews (2018b) (Transpose-Transpose-GEMM-Transpose): for
@@
In his block-scatter matrix tensor contraction (BSMTC) scheme, MatthewsMatthews (2018b) has shown that is possible to hide the transposition cost by loading tensor parts into CPU L2 and L3 cache in order which effectively performs the required permutation of its elements.
Another alternative is the GETT scheme (GEMM-like tensor-tensor multiplication) by Springer and BientinesiSpringer and Bientinesi (2017) which generates code to call matrix-matrix multiplication kernels while again reformulating the sub-matrix-packing. A GPU implementation thereof has become a foundation of the single-node library cuTENSOR.cuT (2019)
@@
Comparing to the software for matrix operations, we currently identify two major practical hurdles for tensor software:
@@
Contrary to BLAS for matrix operations, there is no standardized interface for tensor operations. This causes substantial code duplication.Psarras et al. (2022)
@@
As for standardization, there has been some work in the past, but none of the interfaces have prevailed. BTASShi et al. (2016) has aimed at providing a standard and a basic CPU implementation. TBLISMatthews (2018b) provides BLAS-like tensor calls. Until today, there is no GPU implementation of these. During the 2022 Dagstuhl tensor workshop,Bientinesi et al. (2022) a development of domain-specific tensor language has been initiated, but remains unfinished and undocumented. There has been work-technical specifications of additional aspects of standardization, like tensor-memory distributionValeev (2023); cppreference.com community (2023) and randomized multilinear algebra.Murray et al. (2023) Currently, there are recurring meetings across academia and private sector under the TAPP (Tensor Algebra Processing Primitives) standardization initiative.TAPP working group (2024)
Regarding the implementations, there are numerous scattered efforts listed in Ref. Psarras et al., 2022. However, for open-source distributed memory GPU libraries, there are only a few major players. i) ExaTENSORLyakh (2019); Pototschnig et al. (2021) based on TAL-SHtal (2018), where the key developer has left to industry. ii) TiledArrayPeng et al. (2016) and Cyclops CTF,Solomonik et al. (2013) where the solid support for GPUs is still under development. iii) TACO DISTAL,Yadav, Aiken, and Kjolstad (2022) which has performance issues for higher-order tensors. iv) TAMMMutlu et al. (2023) scales well on supercomputers, but has a potentially problematic dependency of Global Arrays,Palmer et al. (2022) with a small user base for a communication library. v) cuTENSORcuT (2019); Zhang et al. (2020) is proprietary and only supports Nvidia hardware, while vi) hipTensorMillette, Ma, and Karunanidhi (2024) is at an early stage of development. Other industry products like vii) PyTorchPaszke et al. (2019) and viii) TensorFlowAbadi (2016) focus almost exclusively on machine learning and again lack features required in CC.